# Parallel Split Learning with Global Sampling

### v1

@@ IV-C Straggler Mitigation
However, adjusting the $\boldsymbol{\pi}$ can increase the batch deviation, as Corollary 1 would no longer hold. The choice of $\boldsymbol{\pi}$ is thus crucial to balance the trade-off between the batch deviation and the concentration of stragglers in the global batches. To address this issue, we treat $\boldsymbol{\pi}$ as a parameter to be optimized with respect to the desired trade-off. Inspired by the Latent Dirichlet Allocation model [28], we introduce a Dirichlet prior distribution over $\boldsymbol{\pi}$ with concentration parameter $\boldsymbol{\alpha}\in\mathbb{R}^{K}$ and $\alpha_{k}>0$. As a conjugate to the categorical distribution, the Dirichlet distribution is a suitable choice for the prior distribution. The choice of $\boldsymbol{\alpha}$ is crucial to balance the trade-off between batch deviation and total training time. For now we assume that $\boldsymbol{\alpha}$ is given and will discuss its initialization in the next subsection. The prior distribution over $\boldsymbol{\pi}$ is defined as:
@@
Due to the sum inside the logarithm in Equation 2, a closed-form solution for the MAP estimate is not available. Instead, we use the Expectation-Maximization (EM) algorithm [29] to iteratively estimate the mixture proportions $\boldsymbol{\pi}^{*}$.
@@
$=\sum_{n=1}^{N}\sum_{k=1}^{K}\gamma_{nk}\big{\{}\ln\pi_{k}+\ln\beta_{k,y_{n}}\big{\}}$ (3)
Note that we use the expected value of the posterior distribution $P(\boldsymbol{z}\mid\boldsymbol{y},\boldsymbol{\pi}^{\text{old}},\boldsymbol{\beta})$ in Equation 3 since the latent variables $\boldsymbol{z}$ are not observed. The responsibilities $\gamma_{nk}$ are given by:
$\gamma_{nk}=\frac{\pi_{k}^{\text{old}}\beta_{k,y_{n}}}{\sum_{k^{\prime}\in\mathcal{K}}\pi_{k^{\prime}}^{\text{old}}\beta_{k^{\prime},y_{n}}}$
@@
$\boldsymbol{\pi}^{\text{new}}=\operatorname*{arg\,max}_{\boldsymbol{\pi}}\mathcal{Q}(\boldsymbol{\pi},\boldsymbol{\pi}^{\text{old}})+\ln P(\boldsymbol{\pi})$ (4)
@@
Afterwards, $\boldsymbol{\pi}^{\text{old}}$ is updated to $\boldsymbol{\pi}^{\text{new}}$ and the E- and M-steps are iteratively repeated until convergece. The convergence is achieved once $||\boldsymbol{\pi}^{\text{new}}-\boldsymbol{\pi}^{\text{old}}||_{2}<\tau$, where $\tau>0$ is the convergence threshold hyperparameter. Once the algorithm converges, the MAP estimate $\boldsymbol{\pi}^{*}=\boldsymbol{\pi}^{\text{new}}$ is obtained. $\boldsymbol{\pi}^{*}$ is guaranteed to be a local maximum of the posterior distribution $P(\boldsymbol{\pi}\mid\boldsymbol{y},\boldsymbol{\alpha},\boldsymbol{\beta})$. The complete EM algorithm for the MAP estimation is outlined in Algorithm 2.
Note that since $y_{n}=y_{n^{\prime}}\implies\gamma_{nk}=\gamma_{n^{\prime}k}$ for $n,n^{\prime}\in\{1,2,\dots,N\}$, the responsibilities $\gamma_{nk}$ can be precomputed for each class $m\in\mathcal{M}$. We define $\Gamma_{k}\in\mathbb{R}^{M}$ as the classwise responsibilities for client $k$ with
$\Gamma_{k,m}=\frac{\pi_{k}^{\text{old}}\beta_{k,m}}{\sum_{k^{\prime}\in\mathcal{K}}\pi_{k^{\prime}}^{\text{old}}\beta_{k^{\prime},m}}$ (5)
@@
This allows us to reformulate the term $N_{k}$ in the M-step as $N_{k}=\boldsymbol{\nu}^{\top}\Gamma_{k}$. The class counts $\boldsymbol{\nu}$ need to be computed only once at the beginning of the EM algorithm and $\Gamma_{k}$ involves the computation of $M\cdot K$ responsibilities per iteration. In the original formulation, $N\cdot K$ responsibilities need to be computed per iteration. Since usually $M\ll N$, this reformulation can lead to a significant speedup. This means that we can take advantage of the complete dataset $\mathcal{D}_{0}$ by choosing $\boldsymbol{y}=\boldsymbol{y}_{0}$ for a more accurate estimation of $\boldsymbol{{\pi}}^{*}$ without additional computational cost compared to using an $N$-sized sample (except for the computation of $\boldsymbol{\nu}$, which is negligible).
@@
12:         Compute $\Gamma_{k}$ using (5)
13:         $N_{k}\leftarrow\boldsymbol{\nu}^{\top}\Gamma_{k}$triangleright E-Step
@@ IV-D Initialization of Concentration Parameters
$\alpha_{k}\leftarrow\alpha_{k}\cdot\exp\{\Delta\cdot\frac{(\omega_{k}-\bar{\boldsymbol{\omega}})}{s_{\boldsymbol{\omega}}}\}$
$\text{where}\quad\bar{\boldsymbol{\omega}}=\frac{1}{K}\sum_{k=1}^{K}\omega_{k}\quad\text{and}\quad s_{\boldsymbol{\omega}}=\sqrt{\frac{1}{K-1}\sum_{k=1}^{K}(\omega_{k}-\bar{\boldsymbol{\omega}})^{2}}$
In this adjustment, the z-score $\frac{\omega_{k}-\bar{\boldsymbol{\omega}}}{s_{\boldsymbol{\omega}}}$ standardizes the delay times $\omega_{k}$ by measuring how many standard deviations each delay time is from the mean delay time $\bar{\boldsymbol{\omega}}$. By incorporating the z-score into the exponential function, we achieve a smooth adjustment that either increases or decreases $\alpha_{k}$ based on the relative delay time of each client. $\Delta$ is the single hyperparameter that controls the trade-off between the batch deviation and the total training time. It changes the unit of measurement of the z-score in terms of standard deviations. Higher values of $\Delta$ lead to a more aggressive adjustment of the $\alpha_{k}$ and can reduce the total training time, but may also increase the batch deviation and vice versa.
@@ IV-E Latent Dirichlet Sampling (LDS)
We introduce the LDS method that balances the trade-off between batch deviation and total training time. LDS is detailed in Algorithm 3. The procedure is almost identical to UGS, but differs in a few aspects.
First, the client selection probabilities $\boldsymbol{\pi}$ are estimated using the EM algorithm for MAP estimation as described in the previous section. Therefor the trade-off hyperparameter $\Delta$ needs to be provided. The concentration parameter $\boldsymbol{\alpha}$ is initialized as described in the previous section. Note that for $\Delta=0$, the delay times have no impact on $\alpha_{k}$ and the initialization is equivalent to the first step. In this case, the EM algorithm will almost surely converge to the same $\boldsymbol{\pi}$ as in UGS. Therefor, UGS can be seen as a special case of LDS with $\Delta=0$.
Second, the $\boldsymbol{\pi}$ can not be simply renormalized after the dataset of a client $k$ depletes. Instead, the component $k$ is removed from the model and the remaining $\boldsymbol{\pi}$ are re-estimated to ensure that $\boldsymbol{\pi}$ remains a maximum of the posterior distribution. The choice for $\boldsymbol{\pi}$ at the beginning of the re-estimation is a crucial factor for the convergence speed of the EM algorithm. We differentiate between two cases based on the hyperparameter $R\in\{0,1\}$. For $R=1$, we sample $\boldsymbol{\pi}$ from the prior distribution $P(\boldsymbol{\pi}\mid\boldsymbol{\alpha})$ as starting point (re-initialization). For $R=0$, we use the last $\boldsymbol{\pi}$ as starting point (no re-initialization). The choice of $R$ has a significant impact on the convergence speed of the EM algorithm. We provide empirical results in the next section.
@@ V Simulation Results
This section provides a comprehensive evaluation of the proposed UGS and LDS methods in terms of model performance and training time. We compare PSL extended by our sampling methods with several existing frameworks, focusing on metrics such as test accuracy, convergence behavior, and Time-Per-Epoch (TPE). Furthermore, we investigate the effectiveness of LDS for mitigating stragglers, and analyze its time complexity under various conditions.
@@ V-A Model Performance and Training Time
We trained a ResNet18 [38] model on the CIFAR10 [39] dataset using PSL with the proposed UGS and LDS (with $\Delta=0.0$) in the absence of stragglers. We compared the test accuracy with PSL using fixed and proportional local batch sizes like in [22]. For brevity we refer to this default sampling method as Fixed Proportional Local Sampling (FPLS). If the local batch sizes are fixed and identical like in [27], we refer to this sampling method as Fixed Local Sampling (FLS). To ensure that all local batch sizes are integer, we round them to the nearest integer. If the nearest integer is 00, we set the batch size to $1$ to ensure that each client contributes to the training process. We also include other frameworks, namely SL, FL, and SFL, as well as central learning (CL) as a baseline. The hyperparameter choices, hardware setup, and training details can be found in Appendix A.
We split the datasets either in an IID or non-IID manner, while considering different numbers of clients ${K\in\{16,32,64,128\}}$. To achieve a non-IID split, a Dirichlet-based data partition method is used, similar to the Extended Dirichlet Strategy in [16]. The method is applied to the datasets such that each client has exactly $2$ classes with strongly varying dataset sizes, simulating strong heterogeneity among the clients. The resulting client datasets are shown exemplary in Figure 4 for $K=16$ clients. We trained the model for 100 epochs with a global batch size of $B=128$ and report the maximum test accuracy achieved during training in Table II. Accuracies are averaged over $5$ runs with different random seeds and given in percentages. The best accuracy for each split and $K$ is highlighted in boldface. The non-IID split is denoted as $(\overline{\text{IID}})$ in the table. CL achieved a maximum accuracy of 90.25%. In Figure 5, we show the test accuracy per epoch for a run with $K=128$.
@@
To investigate the performance gap between our methods and PSL with fixed local batch sizes, we compared the batch deviation (as defined in Equation 1) of UGS with FPLS and FLS. We calculated the mean and standard deviation of the batch deviation in the global batch sequence generated by the different methods. We considered different values of $B$ and $K$ under IID and non-IID settings, as shown in Figure 6. In the IID scenario, all sampling strategies exhibited nearly identical means and standard deviations. This is consistent with Theorem 1, since $\beta_{0,m}\approx\beta_{k,m}$ for all $k$ and $m$. Differences between UGS, FPLS, and FLS were minimal, with UGS maintaining a nearly constant mean and standard deviation in both IID and non-IID cases. However, FPLS and FLS showed a marked increase in mean and standard deviation in the non-IID scenario. The worst batch deviation in non-IID cases was observed with FLS, while FPLS showed a lower mean and standard deviation that decreased with higher $K$, though still considerably higher than UGS. For UGS, a larger batch size consistently reduced both mean and standard deviation in both IID and non-IID cases, aligning with the deviation bound in Lemma 1. In contrast, for FPLS and FLS, this reduction was observed only in the IID case. This suggests the dominance of the bias term in the deviation bound of Lemma 2 in the non-IID setting. Our findings suggests that sampling methods with fixed local batch sizes introduce greater batch deviation to the training process, explaining the performance gap between UGS and FPLS.
@@ V-B Evaluation of LDS for Straggler Mitigation
To evaluate the effectiveness of the LDS method for straggler mitigation, we introduced stragglers into the training process by defining a straggler probability $p_{s}$, a minimum straggler delay $\omega_{\text{min}}$, and a maximum straggler delay $\omega_{\text{max}}$. We randomly selected a subset of clients to be stragglers with probability $p_{s}$ and introduced a delay $\boldsymbol{\omega}$ uniformly distributed between $\omega_{\text{min}}$ and $\omega_{\text{max}}$. For non-stragglers the delay time was set to zero. Each client $k$ waited for $\omega_{k}$ milliseconds before sending its data to the server. We trained the model again on the CIFAR10 dataset with the same hyperparameters using PSL with LDS. We measured the TPE for different straggler probabilities $p_{s}\in\{0.1,0.2,0.3\}$ and $\Delta\in\{0.5,1.0,1.5\}$ under IID and non-IID dataset splits. The results are shown in LABEL:tab:roundtimes. TPEs are averaged over the $5$ runs and given in seconds. We included the TPE for $\Delta=0.0$ in the absence of stragglers as a baseline. The best TPE for each split, $K$, and $p_{s}$ is highlighted in boldface.
@@
We compared test accuracy for each setting, as shown in Table III. Accuracies are averaged over $5$ runs and given in percentages. The best accuracy for each split, $K$ and $p_{s}$ is highlighted in boldface. The test accuracy remained close to the UGS baseline across all $p_{s}$ and $\Delta$, demonstrating robustness.

### v2

@@ IV-C Straggler Mitigation
However, adjusting the $\boldsymbol{\pi}$ can increase the batch deviation, as Corollary 1 would no longer hold. The choice of $\boldsymbol{\pi}$ is thus crucial to balance the trade-off between the batch deviation and the concentration of stragglers in the global batches. To address this issue, we treat $\boldsymbol{\pi}$ as a parameter to be optimized with respect to the desired trade-off. Inspired by the Latent Dirichlet Allocation model [26], we introduce a Dirichlet prior distribution over $\boldsymbol{\pi}$ with concentration parameter $\boldsymbol{\alpha}\in\mathbb{R}^{K}$ and $\alpha_{k}>0$ and treat $\boldsymbol{z}$ as a latent variable. As a conjugate to the categorical distribution, the Dirichlet distribution is a suitable choice for the prior distribution. The choice of $\boldsymbol{\alpha}$ is crucial to balance the trade-off between batch deviation and total training time. For now, we assume that $\boldsymbol{\alpha}$ is given and will discuss its initialization in the next subsection. The prior distribution over $\boldsymbol{\pi}$ is defined as:
@@
Due to the sum inside the logarithm in Equation 2, a closed-form solution for the MAP estimate is not available. Instead, we use the EM algorithm [27] to iteratively estimate the mixture proportions $\boldsymbol{\pi}^{*}$.
@@
$=\sum_{n=1}^{N}\sum_{k=1}^{K}\gamma_{nk}\big{\{}\ln\pi_{k}+\ln\beta_{k,y_{n}}\big{\}}.$ (3)
Note that we use the expected value of the posterior distribution $P(\boldsymbol{z}\mid\boldsymbol{y},\boldsymbol{\pi}^{\text{old}},\boldsymbol{\beta})$ in Equation 3 since the latent variable $\boldsymbol{z}$ is not observed. The responsibilities $\gamma_{nk}$ are given by:
$\gamma_{nk}=\frac{\pi_{k}^{\text{old}}\beta_{k,y_{n}}}{\sum_{k^{\prime}\in\mathcal{K}}\pi_{k^{\prime}}^{\text{old}}\beta_{k^{\prime},y_{n}}}.$
@@
$\boldsymbol{\pi}^{\text{new}}=\operatorname*{arg\,max}_{\boldsymbol{\pi}}\mathcal{Q}(\boldsymbol{\pi},\boldsymbol{\pi}^{\text{old}})+\ln P(\boldsymbol{\pi}\mid\boldsymbol{\alpha}).$ (4)
@@
Afterwards, $\boldsymbol{\pi}^{\text{old}}$ is updated to $\boldsymbol{\pi}^{\text{new}}$ and the E- and M-steps are iteratively repeated until $||\boldsymbol{\pi}^{\text{new}}-\boldsymbol{\pi}^{\text{old}}||_{2}<\tau$, where $\tau>0$ is the convergence threshold hyperparameter. The complete EM algorithm for the MAP estimation is outlined in Algorithm 2.
Note, that since $y_{n}=y_{n^{\prime}}\implies\gamma_{nk}=\gamma_{n^{\prime}k}$ for $n,n^{\prime}\in\{1,2,\dots,N\}$, the responsibilities $\gamma_{nk}$ can be precomputed for each class $m\in\mathcal{M}$. We define $\hat{\boldsymbol{\gamma}}_{k}\in\mathbb{R}^{M}$ as the classwise responsibilities for client $k$ with
$\hat{\gamma}_{k,m}=\frac{\pi_{k}^{\text{old}}\beta_{k,m}}{\sum_{k^{\prime}\in\mathcal{K}}\pi_{k^{\prime}}^{\text{old}}\beta_{k^{\prime},m}}$ (5)
@@
This allows us to reformulate the term $N_{k}$ in the M-step as $N_{k}=\boldsymbol{\nu}^{\top}\hat{\boldsymbol{\gamma}}_{k}$. The class counts $\boldsymbol{\nu}$ need to be computed only once at the beginning of the EM algorithm and $\hat{\boldsymbol{\gamma}}_{k}$ involves the computation of $M\cdot K$ responsibilities per iteration. In the original formulation, $N\cdot K$ responsibilities need to be computed per iteration. Since usually $M\ll N$, this reformulation can lead to a significant speedup. This means that we can take advantage of the complete dataset $\mathcal{D}_{0}$ by choosing $\boldsymbol{y}=\boldsymbol{y}_{0}$ for a more accurate estimation of $\boldsymbol{{\pi}}^{*}$ without additional computational cost compared to using an $N$-sized sample (except for the computation of $\boldsymbol{\nu}$, which is negligible).
@@
12:         Compute $\hat{\boldsymbol{\gamma}}_{k}$ using (5)
13:         $N_{k}\leftarrow\boldsymbol{\nu}^{\top}\hat{\boldsymbol{\gamma}}_{k}$triangleright E-Step
@@ IV-D Initialization of Concentration Parameters
$\alpha_{k}\leftarrow\alpha_{k}\cdot\exp\{\Delta\cdot\frac{(\omega_{k}-\bar{\omega})}{s_{\boldsymbol{\omega}}}\}$
$\text{where}\quad\bar{\omega}=\frac{1}{K}\sum_{k=1}^{K}\omega_{k}\quad\text{and}\quad s_{\boldsymbol{\omega}}=\sqrt{\frac{1}{K-1}\sum_{k=1}^{K}(\omega_{k}-\bar{\omega})^{2}}$
In this adjustment, the z-score $\frac{\omega_{k}-\bar{\boldsymbol{\omega}}}{s_{\boldsymbol{\omega}}}$ standardizes the delay times $\omega_{k}$ by measuring how many standard deviations each delay time is from the mean delay time $\bar{\omega}$. By incorporating the z-score into the exponential function, we achieve a smooth adjustment that either increases or decreases $\alpha_{k}$ based on the relative delay time of each client. $\Delta$ is the single hyperparameter that controls the trade-off between the batch deviation and the total training time. It changes the unit of measurement of the z-score in terms of standard deviations. Higher values of $\Delta$ lead to a more aggressive adjustment of the $\alpha_{k}$ and can reduce the total training time, but may also increase the batch deviation and vice versa.
@@ IV-E Latent Dirichlet Sampling (LDS)
We introduce a new method, which we call Latent Dirichlet Sampling (LDS), that balances the trade-off between batch deviation and total training time. LDS is detailed in Algorithm 3. The procedure is almost identical to UGS, but differs in a few aspects.
First, the client selection probabilities $\boldsymbol{\pi}$ are estimated using the EM algorithm for MAP estimation as described in the previous section. Therefor, the trade-off hyperparameter $\Delta$ needs to be provided. The concentration parameter $\boldsymbol{\alpha}$ is initialized as described in the previous section. Note that for $\Delta=0$, the delay times have no impact on $\alpha_{k}$ and the initialization is equivalent to the first step. In this case, the EM algorithm will almost surely converge to the same $\boldsymbol{\pi}$ as in UGS. Thus, UGS can be seen as a special case of LDS with $\Delta=0$.
Second, $\boldsymbol{\pi}$ can not be simply renormalized after the dataset of a client $k$ depletes. Instead, the component $k$ is removed from the model and the remaining $\boldsymbol{\pi}$ are re-estimated to ensure that $\boldsymbol{\pi}$ remains a maximum of the posterior distribution. The choice for $\boldsymbol{\pi}$ at the beginning of the re-estimation is a crucial factor for the convergence speed of the EM algorithm. We differentiate between two cases based on the hyperparameter $R\in\{0,1\}$. For $R=1$, we sample $\boldsymbol{\pi}$ from the prior distribution $P(\boldsymbol{\pi}\mid\boldsymbol{\alpha})$ as starting point (re-initialization). For $R=0$, we use the last $\boldsymbol{\pi}$ as starting point (no re-initialization). The choice of $R$ has a significant impact on the convergence speed of the EM algorithm. We provide empirical results in the next section.
@@ V Simulation Results
This section provides a comprehensive evaluation of the proposed UGS and LDS methods in terms of model performance and training time. We compare PSL extended by our sampling methods with several existing frameworks, focusing on metrics such as test accuracy, convergence behavior, and training time per epoch (TPE). Furthermore, we investigate the effectiveness of LDS for mitigating stragglers, and analyze its time complexity under various conditions.
@@ V-A Model Performance and Training Time
We trained a ResNet18 [35] model on the CIFAR10 [36] dataset using PSL with the proposed UGS and LDS (with $\Delta=0.0$) in the absence of stragglers. We compared the test accuracy with PSL using fixed and proportional local batch sizes like in [19]. For brevity we refer to this default sampling method as fixed proportional local sampling (FPLS). If the local batch sizes are fixed and identical like in [24], we refer to this sampling method as fixed local sampling (FLS). To ensure that all local batch sizes are integer, we round them to the nearest integer. If the nearest integer is 00, we set the batch size to $1$ to ensure that each client contributes to the training process. We also include other frameworks, namely SL, FL, and SFL, as well as CL as a baseline. The hyperparameter choices, hardware setup, and training details can be found in Appendix A.
We split the datasets either in an IID or non-IID manner, while considering different numbers of clients ${K\in\{16,32,64,128\}}$. To achieve a non-IID split, a Dirichlet-based data partition strategy is used, similar to the extended Dirichlet strategy in [15]. The strategy is applied to the datasets so that each client has exactly $2$ classes with strongly varying dataset sizes, simulating strong heterogeneity among the clients. The resulting client datasets are shown exemplary in Figure 4 for $K=16$ clients. We trained the model for 100 epochs with a global batch size of $B=128$ and report the maximum test accuracy achieved during training in Table II. Accuracies are averaged over $5$ runs with different random seeds and given in percentages. The best accuracy for each split and $K$ is highlighted in boldface. The non-IID split is denoted as $(\overline{\text{IID}})$ in the table. CL achieved a maximum accuracy of 90.25%. In Figure 5, we show the test accuracy per epoch for a run with $K=128$.
@@
To investigate the performance gap between our methods and PSL with fixed local batch sizes, we compared the batch deviation (as defined in Equation 1) of UGS with FPLS and FLS. We calculated the mean and standard deviation of the batch deviation in the global batch sequence generated by the different methods. We considered different values of $B$ and $K$ under IID and non-IID settings, as shown in Figure 6. For each batch size $B$ we selected 5 different values for $K$ that were evenly distributed in $\{20,\dots,B\}$. In the IID scenario, all sampling strategies exhibited nearly identical means and standard deviations. This is consistent with Theorem 1, since $\beta_{0,m}\approx\beta_{k,m}$ for all $k$ and $m$. Differences between UGS, FPLS, and FLS were minimal, with UGS maintaining a nearly constant mean and standard deviation in both IID and non-IID cases. However, FPLS and FLS showed a marked increase in mean and standard deviation in the non-IID scenario. The worst batch deviation in non-IID cases was observed with FLS, while FPLS showed a lower mean and standard deviation that decreased with higher $K$, though still considerably higher than UGS. For UGS, a larger batch size consistently reduced both mean and standard deviation in both IID and non-IID cases, aligning with the deviation bound in Lemma 1. In contrast, for FPLS and FLS, this reduction was observed only in the IID case. This suggests the dominance of the bias term in the deviation bound of Lemma 2 in the non-IID setting. Our findings suggests that sampling methods with fixed local batch sizes introduce greater batch deviation to the training process, explaining the performance gap between UGS and FPLS.
@@ V-B Evaluation of LDS for Straggler Mitigation
To evaluate the effectiveness of the LDS method for straggler mitigation, we introduced stragglers into the training process by defining a straggler probability $p_{s}$, a minimum straggler delay $\omega_{\text{min}}\in\mathbb{N}$, and a maximum straggler delay $\omega_{\text{max}}\in\mathbb{N}$ with $\omega_{\text{min}}<\omega_{\text{max}}$. The delays are measured in milliseconds. We randomly selected a subset of clients to be stragglers with probability $p_{s}$ and introduced a delay $\boldsymbol{\omega}$ uniformly distributed between $\omega_{\text{min}}$ and $\omega_{\text{max}}$. For non-stragglers the delay time was set to zero. Each client $k$ waited for $\omega_{k}$ milliseconds before sending its data to the server. We trained the model again on the CIFAR10 dataset with the same hyperparameters using PSL with LDS. We measured the TPE for different straggler probabilities $p_{s}\in\{0.1,0.2,0.3\}$ and $\Delta\in\{0.5,1.0,1.5\}$ under IID and non-IID dataset splits. The results are shown in Table IV. TPEs are averaged over the $5$ runs and given in seconds. We included the TPE for $\Delta=0.0$ in the absence of stragglers as a baseline. The best TPE for each split, $K$, and $p_{s}$ is highlighted in boldface.
@@
We compared the test accuracy for each setting, as shown in Table III. Accuracies are averaged over $5$ runs and given in percentages. The best accuracy for each split, $K$ and $p_{s}$ is highlighted in boldface. The test accuracy remained close to the UGS baseline across all $p_{s}$ and $\Delta$, demonstrating robustness.
@@
TABLE IV: Roundtimes for PSL with LDS
Roundtime $p_{s}=0.0$ $p_{s}=0.1$ $p_{s}=0.2$ $p_{s}=0.3$
$K\downarrow/\Delta\rightarrow$ $0.0$ $0.0$ $0.5$ $1.0$ $1.5$ $0.0$ $0.5$ $1.0$ $1.5$ $0.0$ $0.5$ $1.0$ $1.5$
16 9.71 27.13 14.95 12.99 11.94 28.09 19.95 14.82 12.98 33.82 22.33 20.23 18.23
32 11.57 35.93 16.05 14.03 13.68 38.47 23.13 19.20 17.09 39.07 26.24 23.61 19.07
64 13.62 38.29 21.83 18.63 16.19 41.26 26.64 21.07 19.51 39.53 28.56 23.95 23.15
128 15.68 41.55 24.70 20.45 18.11 43.82 28.83 23.94 21.92 44.11 32.26 26.30 24.26